# Diversity-Enhanced Reasoning for Subjective Questions

### v1

@@ Abstract
Large reasoning models (LRM) with long chain-of-thought (CoT) capabilities have shown strong performance on objective tasks, such as math reasoning and coding. However, their effectiveness on subjective questions that may have different responses from different perspectives is still limited by a tendency towards homogeneous reasoning, introduced by the reliance on a single ground truth in supervised fine-tuning and verifiable reward in reinforcement learning. Motivated by the finding that increasing role perspectives consistently improves performance, we propose MultiRole-R1, a diversity-enhanced framework with multiple role perspectives, to improve the accuracy and diversity in subjective reasoning tasks. MultiRole-R1 features an unsupervised data construction pipeline that generates reasoning chains that incorporate diverse role perspectives. We further employ reinforcement learning via Group Relative Policy Optimization (GRPO) with reward shaping, by taking diversity as a reward signal in addition to the verifiable reward. With specially designed reward functions, we successfully promote perspective diversity and lexical diversity, uncovering a positive relation between reasoning diversity and accuracy. Our experiment on six benchmarks demonstrates MultiRole-R1’s effectiveness and generalizability in enhancing both subjective and objective reasoning, showcasing the potential of diversity-enhanced training in LRMs. [^1] The code and data of this work is available at https://github.com/toward-agi/diverse-o1.git.
Diversity-Enhanced Reasoning for Subjective Questions
Yumeng Wangthanks: Equal contribution.    Zhiyuan Fanthanks: Project leader of student team.    Jiayu Liu     Yi R. (May) Fungthanks: Corresponding author Hong Kong University of Science and Technology ywanglu@connect.ust.hk       yrfung@cse.ust.hk
@@ 1 Introduction
Advances in OpenAI o1 and DeepSeek R1 style models Jaech et al. (2024); DeepSeek-AI et al. (2025) with long reasoning chains Wei et al. (2023) have significantly enhanced the models’ capability in various reasoning tasks Yu et al. (2025); Wu et al. (2024b). Extended reasoning chains enable models to systematically divide and conquer complex problems, iteratively refining intermediate results, and methodically verifying the outputs. The DeepSeek R1 style models reached performance comparable to human experts in many objective reasoning tasks such as math Cobbe et al. (2021); Wang et al. (2025a); Guo et al. (2025) and coding Jain et al. (2024). This success highlights that scaling up the test time computation can significantly boost performance on complex, structured tasks relative to standard single-step prompting Jaech et al. (2024); Wei et al. (2023).
Figure 1: A subjective reasoning task example which demonstrates that the perpective shifts may lead to different answers. We illustrate that diverse perspective thinking benefits subjective question answering.
Beyond objective tasks with a single correct answer, there is a growing interest in applying Large Reasoning Models (LRMs) to subjective or open-ended problems Sarıtaş et al. (2025); Guo et al. (2024). This homogenization stems from standard training pipelines which reduce subjective problems to a single ground truth for both supervised learning and reinforcement learning, thereby penalizing diverse perspectives. As illustrated in Figure 1, due to the single inherent thinking process of LRM, the model tends to rely on homogeneous reasoning trajectories without exploring alternative perspectives, hindering their capacity to incorporate diverse reasoning strategies. Prior works’ solutions mainly fall into two categories: (1) Research on multi-agent debates has demonstrated that incorporating diverse role perspectives substantially improves performance on subjective tasks Aoyagui et al. (2025); Cheng et al. (2024); Liu et al. (2025b), but these methods require multiple interacting models; (2) some approaches ensemble solutions from varied decoding paths Wang et al. (2023), yet the stochastic nature of these paths may not necessarily introduce a broader range of perspectives Naik et al. (2024).
To address this gap, we propose to enhance the diversity of the DeepSeek R1 style reasoning chain, by dynamically integrating diverse role perspectives for subjective question answering, including tasks such as moral dilemmas, opinion-based questions, ambiguous question answering, and so on. To begin with, we conduct a pilot study to confirm two key factors: the effectiveness of long reasoning chains in subjective reasoning tasks, and the impact of varying the number of role perspectives incorporated during reasoning. Our findings show that sequential test-time scaling performance improves as reasoning chains are lengthened, but begins to decline once they surpass an optimal length. Similarly, we identify that there exists an optimal range for the number of role perspectives in the reasoning chain. Building upon these insights, we propose a novel framework MultiRole-R1 that enhances the diversity of long chains of thought. Specifically, we optimize the role perspective diversity and the reasoning diversity.
MultiRole-R1 starts with a novel data construction pipeline combining parallel scaling with sequential scaling to effectively incorporate a broader array of role perspectives into the long reasoning chains. We subsequently refine the model through supervised fine-tuning (SFT), instructing the model to learn the multi-role reasoning format and enhance perspective diversity. Furthermore, to enhance the diversity of the reasoning process, we introduce a carefully designed diversity reward function implemented through a multi-role Generalized Reward Policy Optimization (GRPO) framework, taking diversity as an additional reward signal. By training on subjective questions only, MultiRole-R1 boosts the performance of both subjective and objective reasoning tasks. Our contribution can be summarized as follows:
  • We perform a pilot analysis that reveals the scaling law of reasoning length and role perspectives in subjective reasoning.
  • We propose MultiRole-R1, a training paradigm that incorporates unsupervised SFT data generation and GRPO with reward shaping, which effectively promotes perspective and lexical diversity. By training on subjective questions solely, MultiRole-R1 achieves state-of-the-art performance in both subjective and objective reasoning benchmarks.
  • Our experiments and analysis highlight an interesting phenomenon: optimizing for diversity often aligns with, and can even enhance, the accuracy objective. This synergistic relationship underscores our method’s effectiveness and its robust generalizability to other domains.
2 Related Work
Test-Time Scaling
Recent advancements, such as the success of Deepseek-R1 DeepSeek-AI et al. (2025) in demonstrating robust long-chain-of-thought (CoT) reasoning, have brought significant attention to test-time scaling Wang et al. (2025c); Liu et al. (2025a); Li et al. (2025b). Current approaches to test-time scaling can be broadly classified into four categories Zhang et al. (2025). Sequential scaling, as explored by Madaan et al. (2023) and Xiang et al. (2025), iteratively refines a model’s state and output, with each step building on the previous to form a coherent chain of computations. In contrast, parallel scaling Wu et al. (2025); Wang et al. (2024a) leverage parallelism to expand the coverage of reasoning chain, thereby increasing the likelihood of identifying correct solutions. Additionally, hybrid scaling, as proposed in Wang et al. (2024b) and Zhang et al. (2024), combines parallel generation of diverse options with sequential evaluation and refinement to enhance reasoning capabilities. Finally, internal scaling Muennighoff et al. (2025); Ye et al. (2025) empowers models to autonomously allocate computational resources during inference, moving away from reliance on externally guided strategies. While these approaches have demonstrated impressive performance improvements, they primarily focus on enhancing models’ reasoning capabilities in mathematical or coding tasks which contain fixed and verifiable answers. Such methods largely overlook the importance of promoting diversity in reasoning for open-ended questions, which require consideration from multiple perspectives.
Subjective Tasks and LLM Role-Playing
Subjective task is a type of task that differs from objective tasks, such as commonsense reasoning, code generation and arithmetic reasoning Wang et al. (2025b) -which involve questions with definitive correct or incorrect answers. Subjective task, due to its open-ended nature, has answers that may change with perspective-shifts or context change Wang et al. (2025b); Jentzsch and Kersting (2023); Wu et al. (2024a). This includes culture-related question answering Huang and Yang (2023); DURMUS et al. (2024); Huang et al. (2025b), subjective language interpretation Jones et al. (2025), ethical question answering Hendrycks et al. (2021), creative question answering Lu et al. (2024), and so on. LLM role-playing Shao et al. (2023), including multi-role discussion Wang et al. (2024c); Du et al. (2023); Liang et al. (2024) is one of the most used techniques to study subjective task. It features specialized AI systems to simulate assigned personas, ranging from real-life figures to fictional characters Chen et al. (2024); Li et al. (2025a). The capacity to simulate different personas has been shown to elicit human-like responses, and introduce more diverse and creative reasoning paths Naik et al. (2024) when solving LLM-related tasks Wang et al. (2024d). In this work, we adopt a parallel multi-role reasoning at test time to enable the model think from a diverse perspective via independent reasoning paths.
@@ 3 Pilot Analysis
Figure 2: (a) The performance of Deepseek-r1-distill-qwen-7b DeepSeek-AI et al. (2025) under different reasoning length settings across different datasets. The bar chart shows that longer reasoning chains result in higher accuracy on subjective tasks.  (b) The performance of Deepseek-r1-distill-qwen-7b DeepSeek-AI et al. (2025) under different reasoning length settings across different datasets.  (c) Demonstration of the number of distinct opinions increases as more roles are involved in a single reasoning chain.
Drawing on evidence that diverse perspectives enhance performance on subjective tasks Muscato et al. (2025); Hayati et al. (2024), we introduce perspective diversity into Deepseek R1-style reasoning chains via test-time scaling Snell et al. (2024). We employ two complementary strategies: sequential scaling to deepen reasoning through iterative reflection Zeng et al. (2025), and parallel scaling to enhance exploration by generating multiple reasoning paths Liu et al. (2025c); Rodionov et al. (2025). We evaluate this approach on three subjective reasoning datasets: GlobalOpinion QA DURMUS et al. (2024), BBQ Parrish et al. (2022), and ETHICS Hendrycks et al. (2021) across four decoding strategies adapted from Jiang et al. (2025), with examples provided in Appendix 7:
As shown in Figure 2 (a), increasing the amount of test-time reasoning through scaling indeed yields performance gains. We thus incorporate More think as our baseline.
@@
Having established the effectiveness of test-time scaling, a natural follow-up question is whether longer reasoning chains always result in better performance. However, prior studies on mathematical and commonsense reasoning have shown that excessively long reasoning chains do not necessarily lead to improved outcomes Sui et al. (2025); Ballon et al. (2025); Qin et al. (2025). Therefore, to control the reasoning length for sequential scaling leveraging concepts from budget forcing Muennighoff et al. (2025); Huang et al. (2025a), we conduct an investigation experiment aimed at roughly estimating the optimal reasoning length specifically for subjective problems. We adhere to the experimental setup proposed by Jiang et al. (2025) and experiment on the optimal number of replacing the end-of-thinking delimiter. As shown in Figure 2 (b), increasing the number of "Wait" tokens generally leads to performance improvements across most tasks. For GLOQA, performance continues to improve with more “Wait” insertions, whereas for BBQ and ETHICS, the gains peak around three insertions and diminish or even degrade beyond that point. Based on this observation, we adopt three “Wait” tokens as a balanced setting that achieves clear performance gains without excessive computational overhead. We refer to this configuration as the More think strategy, which not only fosters more cautious and reflective reasoning, but also enables longer reasoning chains capable of supporting more nuanced and diverse role-based perspectives.
@@
We explore the relationship between the number of roles and the answer diversity and consensus accuracy on subjective QA tasks. Intuitively, each additional role contributes a new viewpoint, expanding the solution space. Lu et al. (2024) shows multi-agent debates with 3 fixed roles achieve optimal creativity, but scaling different role perspectives in a single reasoning chain at test time remains unexplored. We conduct a pilot analysis that directly instructs the model to think from $n$ different viewpoints, with $n$ from 1 to 6. Results in Figure 2 (c) show that the number of distinct opinions increases as more roles are involved. In general, we observed that the number of distinct opinions plateaus as $n=4$. Hence, in the following study, we control the number of generated roles $n=2,3,4$ for efficient generation.
Figure 3: Illustration of our MultiRole-R1framework. Stage 1: we collect our data by utilizing budget forcing (sequential scaling) and contrastive role reasoning, sampling N outputs from the decoder (parallel scaling). Stage 2: we automatically construct our training data that integrates various perspectives into a single reasoning chain, and then fine-tune the model to follow the multi-role reasoning format. Stage 3: we utilize GRPO with reward shaping to optimize the reasoning accuracy and diversity. Two accuracy evaluation approaches are applied, depending on whether the ground-truth varies by roles. By taking diversity as an additional reward, we encourage exploration of the reasoning chain, expanding the answer search space.
4 Method
As illustrated in Figure 3, our framework consists of three stages: parallel multi-role reasoning, multi-role finetuning and multi-role reinforcement learning. Formally, given an input question $\mathcal{Q}$ and a reasoning model $\mathcal{M}$, our goal is to diversify the reasoning path $\mathcal{T}$.
4.1 Parallel Multi-Role Reasoning
Multi-Role Exploration
To model multi-perspective reasoning, we first identify a set of $n$ context relevant roles (e.g., domain experts, stakeholders, or personas) through few-shot prompting Brown et al. (2020), denoted as =$\mathcal{R}=\{R_{1},R_{2},...,R_{n}\}$. In particular, we prompt the model to generate roles that may have conflicting viewpoints. The motivation for this is to explore diverse available perspectives. Given candidate roles $\mathcal{R}$, we define the selection probabilities:
$\begin{split}P(R_{i}|\mathcal{Q})=softmax(\mathbb{E}[\mathcal{M}(R_{i}|\mathcal{Q})]\\ +\lambda\mathbb{E}_{R_{i}}[1-\text{sim}(R_{i},R_{j})]),\end{split}$ (1)
where $sim(R_{i},R_{j})=cos(\mathbf{h}_{R_{i}},\mathbf{h}_{R_{j}}|\mathcal{Q})$ and $\mathbf{h}$ denotes the embedding of the LLM $\mathcal{M}$. These terms represent that, in the role selection process, we prioritize role answers that are relevant to the question $\mathcal{Q}$ and contrastive to the existing opinions.
@@ 4.1 Parallel Multi-Role Reasoning
For each role $R_{i}$, we generate $k$ candidate reasoning paths $\mathcal{M}(Q,R_{i})=\mathcal{T}_{R_{i}}=\{T_{R_{i}}^{(1)},T_{R_{i}}^{(2)},...,T_{R_{i}}^{(k)}\}$. This is achieved by adjusting the sampling temperature $\tau=1$ to encourage diverse reasoning paths. To ensure the coherence among different responses of each role, we then apply self-consistency filtering Chen et al. (2023); Wang et al. (2025d) through majority voting and only keep the most consistent answer:
$\hat{T}_{R_{i}}=argmax\sum_{j=1,T\in\mathcal{T}_{R_{i}}}^{k}\mathbbm{1}(T\equiv T_{R_{i}}^{(j)}),$ (2)
@@
4.2 Multi-Role Finetuning
To enhance our model’s reasoning capabilities in both accuracy and diversity, we first fine-tune the base model with self-consistency filtered data. This initial stage teaches the model to adopt a multi-role reasoning format, which involves shifting perspectives throughout its reasoning process. To further promote this diversity, our supervised fine-tuning (SFT) dataset includes 968 examples specifically designed for varied role generation.
Sequential Multi-Role Merging
Given $m$ filtered role perspectives $\{\hat{T}_{R_{1}},...,\hat{T}_{R_{m}}\}$, we generate random combination of role orderings $\Pi$ to avoid the effect of position bias Zheng et al. (2023a). For example, given a multi-role combination $\pi=\{R_{i},R_{j},R_{k}\}$, we construct the training data as follows:
$\mathcal{D}_{\text{train}}=\bigcup_{\pi\in\Pi}\{(\mathcal{Q}\oplus\hat{T}_{R_{i}}\oplus\hat{T}_{R_{j}}\oplus\hat{T}_{R_{k}})\mid\pi\}.$ (3)
We consider two merging strategies depending on task type to allow dynamic integration of role reasoning traces while respecting the nature of the reasoning task: (1) Divergent merging: when the roles are supposed to give different answers, we encourage diverse perspectives to maximize coverage of the searching space. Final prediction is derived via weighted aggregation over differing viewpoints. (2) Convergent merging: when the roles are supposed to give consistent answers, they reach consensus via majority voting within a single sequence, which is in nature soft majority voting within a single sequence.
@@ 4.2 Multi-Role Finetuning
To ensure the quality of the SFT data, we apply several filtering strategies to the merged dataset. To mitigate verbosity bias Zheng et al. (2023b) and reasoning shortcut behaviors, we remove entries in the top and bottom 10th percentiles of answer length. We also discard instances with formatting errors or invalid string patterns. This results in the final 2,700 SFT training data. Additionally, since the gold labels are accessible in the original dataset, we apply a supervised, ground-truth-guided filtering approach as a comparison to our unsupervised self-consistency method. Results of this comparison are presented in Section 7.1.
4.3 Multi-Role Reinforcement Learning
We adopt Group Relative Policy Optimization (GRPO) Shao et al. (2024) for Multi-Role reinforcement learning, which is trained on top of the SFT model. GRPO optimizes the policy by sampling a group of candidate outputs for each prompt and comparing their reward. We incorporate two types of rewards: an accuracy reward $R_{\text{acc}}$ provided by a verifiable reward model that checks answer correctness, and a diversity reward $R_{\text{div}}$ computed from the input text as a shaping signal. The total shaped reward is given by $R=0.9*R_{\text{acc}}+0.1*R_{\text{div}}$. This follows the reward-shaping paradigm Ng et al. (1999), where the auxiliary $R_{\text{div}}$ guides learning without changing the optimal policy.
During training, we observe a synergetic effect of optimizing the diversity and accuracy objectives. This also mitigates issues observed in the SFT baseline, such as excessive verbosity and repetitive reasoning Toshniwal et al. (2025). Finally, note that GRPO computes group advantages $A_{1},A_{2},\dots,A_{G}$ instead of standard reward, which is given by:
$\hat{A}_{i,t}=(R_{i,t}-\mu)/\sigma,t\in\{1,\dots,|G|\}.$
Hence, a group with uniform rewards (all 0s or all 1s) would give zero advantage and stall learning. By adding the diversity term, we ensure intra-group reward variance, enabling informative gradients and continued optimization.
5 Experiment
Model BBQ GLOQA CALI (O.O.D) ETHICS CSQA (O.O.D) GSM8K (O.O.D)

### v3

@@ Abstract
Large Reasoning Models (LRMs) with long chain-of-thought capabilities, optimized via reinforcement learning with verifiable rewards (RLVR), excel at objective reasoning tasks like mathematical problem solving and code generation. However, RLVR is known for degrading generation diversity, which causes LRMs to fall short on subjective reasoning that has multiple answers depending on different role perspectives. While recent studies recognize the importance of diversity-enhanced training in objective reasoning, limited attention has been given to subjective tasks. In this paper, we find that subjective reasoning can be improved by introducing perspective diversity and token-level diversity, with the former one providing a coherent scaffolding anchored to a real-world stakeholder group and the latter one broadening the answer search space. We propose MultiRole-R1, a diversity-enhanced training framework featuring an unsupervised data construction pipeline that synthesizes reasoning chains incorporating various role perspectives. It also employs reinforcement learning via Group Relative Policy Optimization with reward shaping, taking diversity as a reward signal in addition to verifiable reward. Training on subjective tasks solely, MultiRole-R1 increases the in-domain and out-of-domain accuracy by 14.1% and 7.64%, and even enhances the performance on advanced math reasoning such as AIME 2024. We further show that diversity is a more consistent indicator of accuracy than reasoning length. The code and data of this work is available at https://github.com/toward-agi/diverse-o1.git.
@@ 1 Introduction
Advances in DeepSeek-R1 (DeepSeek-AI et al., 2025) and OpenAI o1-style (Jaech et al., 2024) models with long Chain-of-Thoughts (CoT) capabilities (Wei et al., 2023) have substantially improved performance on challenging reasoning tasks, particularly in objective domains such as commonsense (Talmor et al., 2019) and mathematical reasoning (Yu et al., 2025; Wu et al., 2024b; Cobbe et al., 2021; Wang et al., 2025a; Guo et al., 2025, inter alia). Notably, this type of model is trained via reinforcement learning with verifiable rewards (RLVR), which induces a diversity degradation in the model generation (Song et al., 2025b; Dang et al., 2025; Zhao et al., 2025; Wu et al., 2025). This greatly undermines the real-world application, since diversity is crucial for effective sampling for test-time scaling (Yue et al., 2025). Recent studies offer several solutions to enhance diversity in RL training (Song et al., 2025a; Yan et al., 2025), but they mostly focus on objective reasoning.
In contrast to objective tasks, subjective questions are fundamentally different from objective questions since there are no definitive right or wrong answers (Khurana et al., 2024; van der Meer et al., 2024; Wang et al., 2025b; Jentzsch & Kersting, 2023; Wu et al., 2024a): the responses can vary greatly depending on the role or stakeholder perspective. This challenge cannot be solved by current diversity-enhanced training approaches in objective domain since they rely on a single ground truth in optimization. This design inherently trains the model to find one correct answer, making it incapable of generating reasoning that arrives at the multiple valid outcomes required by subjective questions. To tackle this, existing research on subjective reasoning mainly falls in two categories: multi-agent debate (Aoyagui et al., 2025; Cheng et al., 2024; Liu et al., 2025b) and prompting-based methods (Wang et al., 2024b; Lv et al., 2024), with no training methods specifically designed for subjective questions.
In this paper, we address this challenge by proposing MultiRole-R1, a diversity-enhanced RL training framework to improve LLM subjective reasoning. Specifically, we incorporate two levels of diversity: (1) Semantic-level diversity (or perspective diversity), which trains the model to incorporate multiple relevant real-world stakeholder perspectives; (2) Token-level diversity, which broadens the search space of the reasoning chains. In particular, we argue that role perspective diversity is key to this challenge: instead of just seeking random variation, roles provide coherent scaffolding that ensures the diverse outputs are semantically relevant and anchored to real-world groups and stakeholders’ viewpoints (Xu et al., 2025; Wang et al., 2025b).
Figure 1: Illustration of MultiRole-R1 framework. Stage 1 (enhance perspective diversity): LRMs generate seed roles with contrastive opinions, and sample diverse reasoning paths from different roles. We concatenate paths from different perspectives into a CoT, and then finetune the model to follow the multi-role reasoning format. Stage 2 (enhance token-level diversity): we utilize GRPO with diversity reward shaping. Verifiable rewards are applied depending on whether the ground-truth varies by roles. We take diversity as an additional reward to promote exploration efficiency.
We conduct a pilot analysis to determine the optimal number of roles and the reasoning length of the generated paths. MultiRole-R1 subsequently finetunes the model on self-synthesized reasoning paths to enhance semantic-level diversity, instructing it to self-teach on multi-role generation and role reasoning, as shown in Figure 1. Furthermore, to enhance the token-level diversity, MultiRole-R1 applies a diversity reward function combining an array of existing token-level diversity metrics, such as lexical diversity, structural diversity, and discourse diversity. This is used as a reward signal in addition to the verifiable reward in Generalized Reward Policy Optimization (GRPO).
To evaluate the effectiveness and generalizability of our approach, we train DeepSeek-R1 series models and Qwen-3-8B using MultiRole-R1 and test them on both subjective and objective questions. Results show that MultiRole-R1 boosts performance by an average of 14.1% on three in-domain (ID) subjective tasks, and 7.64% on four out-of-domain (OOD) tasks that include both subjective and objective questions. Interestingly, our approach even achieves a performance gain on the OOD advanced math reasoning dataset AIME 2024 by 5.78%. Our further analysis shows that among the 10.6% average performance gain of MultiRole-R1, 8.3% is contributed by multi-role SFT, and 6.6% is contributed by GRPO with diversity reward shaping. This verifies the necessity of incorporating perspective diversity in subjective reasoning, and also corroborates the cruciality of token-level diversity in test-time scaling. Moreover, we find a strong per-task correlation between diversity and accuracy ($r=0.74$), which markedly outweighs the correlation between length and accuracy ($r=0.55$). This result extends the previous finding of a correlation between diversity and task performance in objective tasks to subjective tasks, indicating that diversity is a more consistent indicator of accuracy than reasoning length. Our contributions can be summarized as follows:
  • To our knowledge, we are the first to introduce diversity-enhanced training for subjective reasoning tasks. We propose MultiRole-R1, a training paradigm that incorporates unsupervised reasoning path synthesis and GRPO with diversity reward shaping, which effectively enables LRM to include diverse perspectives and generate multiple different answers in subjective reasoning.
  • We verify MultiRole-R1 on four models using solely subjective questions for training. Results show that the models achieve state-of-the-art performance in three ID and four OOD tasks, and can generalize to advanced math reasoning such as AIME 2024.
  • Our analysis highlights diversity as a more consistent indicator of accuracy than reasoning length.
@@ 2 Pilot Analysis
Figure 2: (a) The performance of Deepseek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025) under different reasoning length settings across different datasets. The bar chart shows that longer reasoning chains result in higher accuracy on subjective tasks.  (b) Accuracy gain of Deepseek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025) when trailing with more wait tokens.  (c) Demonstration of the number of distinct opinions increases as more roles are involved in a single reasoning chain.
Inspired by previous work (Zelikman et al., 2022; Peng et al., 2025) that fine-tuning on self-synthesized CoT improves reasoning, we extend this methodology by enabling the model to self-improve on synthesizing diverse perspectives into a single reasoning chain. To achieve this, we concatenate multiple single-role reasoning chains into one long reasoning path, as shown in Figure 1 Stage 1. One natural question is how to decide the format of the multi-role reasoning chain: how many role perspectives to include, and how long should the reasoning path be? On the one hand, we want to incorporate more roles to cover comprehensive perspectives and leverage the self-correction ability of the long reasoning chain. On the other hand, we want to control the reasoning length within an optimal range, as excessive verbosity leads to performance degradation (Hassid et al., 2025) and is computationally expensive.
To answer the question, our pilot analysis aims to study the optimal number of roles and reasoning length. We primarily consider Deepseek-R1-Distill-Qwen-7B, and a set of subjective questions sampled from GlobalOpinion QA (DURMUS et al., 2024) and BBQ (Parrish et al., 2022) datasets. To elicit multiple viewpoints in a single reasoning chain, we employed budget-forcing (Muennighoff et al., 2025), which replaces the end-of-thinking token with a continuation token (i.e., “wait, I need to think from {role}’s perspective”), to divert the model to a different role perspective. These roles, designed to be mutually contrasting in opinions, were pre-generated by the base model via prompting.
Besides budget-forcing that produces longer paths than the regular reasoning, we also compare it with settings that are shorter than regular reasoning, with examples in Appendix C.2. We observe that reasoning lengths longer than regular think (i.e., more think) significantly outperform other settings, as shown in Figure 2 (a). This motivates us to focus on more think setting, which is about how many wait tokens should be appended. The following highlights the main findings of the pilot analysis:
@@
As shown in Figure 2 (b), increasing the number of “Wait” tokens generally leads to performance improvements across most tasks, where the gains mostly peak around three insertions and diminish or even degrade beyond that point.
@@
Results in Figure 2 (c) show that the number of distinct opinions increases as more roles are involved, with number of roles $n=3$ as a turning point where the increase of roles provides less salient information gain compared to the previous. Hence, we incorporate three roles in the path generation.
3 Methodology
As illustrated in Figure 1, our framework consists of two stages: multi-role reasoning paths synthesis & finetuning and diversity-enhanced reinforcement learning. Formally, given an input subjective question $\mathcal{Q}$ and a reasoning model $\mathcal{M}$, our goal is to diversify the reasoning path $\mathcal{T}$.
3.1 Multi-Role Reasoning Paths Synthesis & Finetuning
The objective of this stage is to enhance perspective diversity: besides training the model to “think deeper from its perspective” (DeepSeek-AI et al., 2025; Jaech et al., 2024), we also train the model to consider “from which perspective to think”.
Multi-Role Exploration and Sampling
To model multi-perspective reasoning, we first identify $n$ context-relevant roles (e.g., domain experts, stakeholders, or personas) through few-shot prompting (Brown et al., 2020), denoted as $\mathcal{R}=\{\mathcal{R}_{1},\mathcal{R}_{2},...,\mathcal{R}_{n}\}$. In particular, we prompt the model to generate roles with conflicting viewpoints. The motivation for this is to explore diverse available perspectives. Given a candidate role $\mathcal{R}_{i}$, LLM $\mathcal{M}$ and question $\mathcal{Q}$, we define the selection probabilities:
$\begin{split}P(\mathcal{R}_{i}|\mathcal{Q})=softmax(\mathbb{E}[\mathcal{M}(\mathcal{R}_{i}|\mathcal{Q})]+\alpha\mathbb{E}_{\mathcal{R}_{i}}[1-\text{sim}(\mathcal{R}_{i},\mathcal{R}_{j})]),\end{split}$ (1)
where $sim(\mathcal{R}_{i},\mathcal{R}_{j})=cos(\mathbf{h}_{\mathcal{R}_{i}},\mathbf{h}_{\mathcal{R}_{j}}|\mathcal{Q})$ and $\mathbf{h}$ denotes the LLM embedding. The intuition for this is to prioritize role answers that are relevant to the question and contrastive to the existing opinions.
@@ 3.1 Multi-Role Reasoning Paths Synthesis & Finetuning
For each role $\mathcal{R}_{i}$, we sample $k$ reasoning paths from the decoder with temperature $\tau=1$, denoted as $\mathcal{M}(Q,\mathcal{R}_{i})=\mathcal{T}_{\mathcal{R}_{i}}=\{\mathcal{T}_{\mathcal{R}_{i}}^{(1)},\mathcal{T}_{\mathcal{R}_{i}}^{(2)},...,\mathcal{T}_{\mathcal{R}_{i}}^{(k)}\}$. To ensure the coherence among different responses of each role, we then apply self-consistency filtering (Chen et al., 2023; Wang et al., 2025c) through majority voting and only keep the most consistent answer:
$\hat{\mathcal{T}}_{\mathcal{R}_{i}}=argmax\sum_{j=1,\mathcal{T}\in\mathcal{T}_{\mathcal{R}_{i}}}^{k}\mathbbm{1}(\mathcal{T}\equiv\mathcal{T}_{\mathcal{R}_{i}}^{(j)}),\vskip-7.5pt$ (2)
@@
Reasoning Structure Generation
Given $m$ filtered role perspectives $\{\hat{\mathcal{T}}_{\mathcal{R}_{1}},...,\hat{\mathcal{T}}_{\mathcal{R}_{m}}\}$, we generate random combinations of role orderings $\Pi$ to avoid the effect of position bias (Zheng et al., 2023a). For example, given a multi-role combination $\pi=\{\mathcal{R}_{i},\mathcal{R}_{j},\mathcal{R}_{k}\}$, we construct the training data as:
$\mathcal{D}_{\text{train}}=\bigcup_{\pi\in\Pi}\{(\mathcal{Q}\oplus\hat{\mathcal{T}}_{\mathcal{R}_{i}}\oplus\hat{\mathcal{T}}_{\mathcal{R}_{j}}\oplus\hat{\mathcal{T}}_{\mathcal{R}_{k}})\mid\pi\}.\vskip-7.5pt$ (3)
We consider two merging strategies depending on task type to allow dynamic integration of role reasoning paths: (1) divergent merging: for tasks where roles are expected to provide different answers, the final prediction is derived through a weighted aggregation of the various viewpoints; (2) convergent merging: for tasks where roles should yield a consistent answer, a consensus is reached via majority voting within the reasoning sequence.
@@
To ensure data quality, we apply several filtering strategies to the merged data. To mitigate verbosity bias (Zheng et al., 2023b) and reasoning shortcut behavior, we remove the top and bottom 10th percentiles of responses by length. We also discard instances with formatting errors or invalid string patterns. This filtering process yielded a final training set of 2,700 entries. For comparison against our self-consistency filtering method, we also applied a supervised ground-truth filtering approach. Comparison results of two filtering strategies are in Section 5.
3.2 Diversity Enhanced Reinforcement Learning
This stage aims to enhance the diversity of the reasoning chain, broadening the answer search space. We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) for Multi-Role reinforcement learning, which is trained on top of the SFT model. GRPO optimizes the policy by sampling a group of candidate outputs for each prompt and comparing their reward. We incorporate two types of rewards: an accuracy reward $\textbf{R}_{\text{acc}}$ provided by a verifiable reward model that checks answer correctness, and a diversity reward $\textbf{R}_{\text{div}}$ computed from the input text as a shaping signal. The total shaped reward is formulated by $\textbf{R}=\delta\textbf{R}_{\text{acc}}+(1-\delta)\textbf{R}_{\text{div}}.$
This follows the reward-shaping paradigm (Ng et al., 1999), where the auxiliary $\textbf{R}_{\text{div}}$ guides learning without changing the optimal policy. Detailed setting is presented in Appendix B.
During training, we observe a synergetic effect of optimizing the diversity and accuracy objectives. This also mitigates issues observed in the SFT baseline, such as excessive verbosity and repetitive reasoning (Toshniwal et al., 2025). Finally, note that GRPO computes group (G) advantages $\textbf{A}_{1},\textbf{A}_{2},\dots,\textbf{A}_{G}$ instead of standard reward, which is given by: $\textbf{A}_{i}=(\textbf{R}_{i,t}-\mu)/\sigma,t\in\{1,\dots,|G|\}.$ Hence, a group with uniform rewards (all 0s or all 1s) would give zero advantage and stall learning. By adding the diversity term, we ensure intra-group reward variance, enabling informative gradients and continued optimization. Mathematical proofs and detailed derivations are provided in Appendix A.
4 Experiment and Results
Table 1: Main results of the baselines (specified in Section 4.2) and our proposed method. Acc. is the pass@1 accuracy of the task (in %) and Div. measures the length normalized diversity score of the reasoning chain (in %). We include two ablations of MultiRole-R1, including SFT on self-consistency filtered data only (Ours SelfConsis SFT), and also SFT with vanilla GRPO (Ours SelfConsis SFT + GRPO). “GRPO(RS)” represents GRPO with reward shaping, which is used in MultiRole-R1. OOD denotes the datasets that are for testing only.
Model BBQ GLOQA ETHICS CALI (OOD) CSQA (OOD) GSM8K (OOD)